Calculate the energy required to excite a particle in a cubic box of edge length 1Å from its ground state to its first excited state.
1
Expert's answer
2020-11-02T08:49:01-0500

E1= h^2/8mL^2 = (6.63x10^-34)^2/8x9.1x10^-31 x 10^-10 x 10^-10 = 0.6038 x 10^-17J.